OpenStudy (anonymous):

Multiply. State any excluded values. q+5/-3 * -3q/q+2

OpenStudy (anonymous):

\[\frac{q+5}{-3}\times \frac{-3q}{q+2}\]?

OpenStudy (anonymous):

only thing i can see to do is cancel a \(-3\) top and bottom and get \[\frac{(q+5)q}{q+2}\]

OpenStudy (anonymous):

there has to be another step because the answers i have are multiple choice and thats not one of them.

OpenStudy (anonymous):

maybe \[\frac{q^2+5q}{q+2}\]?

OpenStudy (anonymous):

also not one of the options. would you like me to list the options and see if you can figure out what to do by looking at the answers?

OpenStudy (anonymous):

did i interpret the question correctly?